Use the van der Waals equation of state to calculate the pressure of 3.90 mol of H20 at 489 K in a 4.20 L vessel. Van der Waals constants can be found here. Number atm Use the ideal gas equation to calculate the pressure under the same conditions. Number atm
Answer #1

P   =

V = 4.2L

T = 489K

n = 3.90 moles

a   = 5.46atm L^2/mole^2

b   = 0.03049L/mole

( P + aV^2/n^2)(v-nb) = nRT

(P + 5.46*(4.2)^2/(3.9)^2 )(3.9-4.2*0.03049) = 3.9*0.0821*489

( P + 6.3323)(3.771942) = 156.57291

P   = 35.17atm

b.

PV = nRT

P   = nRT/V

       = 3.9*0.0821*489/4.2

        = 37.28atm